OpenStudy (anonymous):

find the derivative x^1/4 (written as a square root), x=625 b) Find the equation of the tangent line at x=625

OpenStudy (anonymous):

For my derivative, i had 1/4x^3/4 when i sub in 625, i get 1/500 is that right?

OpenStudy (anonymous):

trying to get my y-value,

OpenStudy (anonymous):

Lol...Didnt Get...u Find the Derivative of ---- ?

OpenStudy (anonymous):

\[\sqrt[4]{x}\]

OpenStudy (anonymous):

is what i need the derivative of, then i need to sub in 625 into the derivative to get x.

OpenStudy (anonymous):

\[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\] \[\sqrt[4]{x} = x ^{\frac{ 1 }{ 2 }*\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

\[\frac{ d }{ dx }(x ^{\frac{ 1 }{ 8 }})\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ 8 }* x ^{\frac{ 1 }{ 8 }-1}\]

OpenStudy (anonymous):

\[\sqrt[4]{x} = x^{1/4}\]

OpenStudy (anonymous):

not a half times a quarter.

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 }*x ^{\frac{ -7 }{ 8 }} = \frac{ 1 }{ 8 }*\frac{ 1 }{ x ^{\frac{ 7 }{ 8 }} }\]

OpenStudy (anonymous):

\[625 = 5^4\]

OpenStudy (anonymous):

Hope nw u can solve...)

OpenStudy (anonymous):

i dont get what you're doing here.

OpenStudy (anonymous):

i got the derivative - \[\frac{ 1 }{ 4 }x ^{\frac{ -1 }{ 4 }}\]

OpenStudy (anonymous):

and when i sub in 625, i get 1/500 as my x value for the equation.

OpenStudy (anonymous):

that is not the \(x\) value, that is the slope

OpenStudy (anonymous):

actually \[\frac{1}{4\sqrt[4]{625}}=\frac{1}{4\times 2}=\frac{1}{20}\]

OpenStudy (anonymous):

ooops i mean \[\frac{1}{4\times 5}=\frac{1}{20}\]

OpenStudy (anonymous):

oops still wrong derivative is \[\frac{1}{4\sqrt[4]{x^3}}\] so you are correct, you get \(\frac{1}{500}\) as your slope

OpenStudy (anonymous):

I don't think you guys are understanding my question. I need to find the derivative of the given equation, then i have to find d/dx (625) so i have to sub in 625 into my derivative, so that gave me 1/500 for my slope. so i then sub in 1/500 into the derivative equation to get my x value, then i'll sub that x value into my original equation to get my y value.

OpenStudy (anonymous):

ok you have the function \[f(x)=\sqrt[4]{x}\] now \(f(625)=5\) so you know the point on the graph is \((625,5)\) the slope is \(f'(625)=\frac{1}{500}\) to find the equation of the line tangent to the curve at \((625,5)\) use the point slope formula and get \[y-5=\frac{1}{500}(x-625)\] you can clean that up with some algebra if you like, depending on what you need to do with it

OpenStudy (anonymous):

that is the answer to the question " Find the equation of the tangent line at x=625"

OpenStudy (anonymous):

thank you, sorry about the late delay... open study stopped working for me.